OpenStudy (anonymous):

Solve for R in the formula V = IR, when I = 5 and V = 15.

OpenStudy (anonymous):

try to isolate R by either dividing or multiplying then replace the variables and solve

OpenStudy (anonymous):

R = V/I = 15/5 = 3

OpenStudy (anonymous):

dont give them the answers they have to find them by themselves but you can help them out

OpenStudy (anonymous):

V=IR can be converted into: R=V/I substitute in the values given for V and I and solve for R. R=15/5 R=3

OpenStudy (anonymous):

luls

OpenStudy (anonymous):

@Rohith15 @bheaze dont give them the answers they have to find them by themselves we are here to help not do their homework

OpenStudy (anonymous):

i reckon this site's traffic would more than double if answers are provided